OpenStudy (lncognlto):

Solve the equation sqr3 sinx = 1/2 sec x for 0 deg < x < 360 deg.

OpenStudy (foreverandalways13):

the magic of a calculator...lol

OpenStudy (foreverandalways13):

ok nvmd....idk I have noooo idea what ur doing kid....

OpenStudy (lncognlto):

lol @leah

OpenStudy (jasmineflvs):

what kind of problem is it??

OpenStudy (jdoe0001):

$$\bf \large { \sqrt{3} sin(x) = \frac{1}{2}sec(x)\\ \sqrt{3} sin(x) = \frac{1}{2cos(x)}\\ sin(x)cos(x) = \frac{1}{2}\frac{1}{\sqrt{3}} \implies \frac{a}{c}\frac{b}{c}\\ sin(x) = \frac{1}{2}\\ cos(x) = \frac{1}{\sqrt{3}} } $$

OpenStudy (jdoe0001):

or the other way around

OpenStudy (jdoe0001):

then just arcCosine and arcSine both sides to get the angles

OpenStudy (jdoe0001):

\(\bf sin(x)cos(x) = \frac{1}{2}\frac{1}{\sqrt{3}} \implies \frac{b}{c}\frac{a}{c}\) had them backward :/

OpenStudy (jasmineflvs):

what kind of problem

OpenStudy (lncognlto):

Okay, so sin x = 1/sqr3 and cos x = 1/2?

OpenStudy (jdoe0001):

or the other way around, either way will give 1/2 secx

OpenStudy (jdoe0001):

if you recall your 45-45-90 rule and your 30-60-90 rule you'd know who those guys are :)

OpenStudy (lncognlto):

Okay, thanks.

OpenStudy (jdoe0001):

yw